A rectangular metal block is 0.6 m long, 0.2 m wide, and 0.1 m tall. The block weighs 270 N. The block is lowered into an unknown fluid until it is completely submerged. In the fluid, the block weighs 164 N. What is the density of the fluid (in kg/m3)? Use g = 9.79 m/s2.
Answer #1

volume of block = 0.6 * 0.2 * 0.1 = 0.012 m^3 = vol of liq displaced

w = 270 N

when lowered in liquid,

w = 164N

Buyonce force,B = 270 - 164 = 106N

Now,

   \(B = \rho vg\)   

   \(106 = \rho * 0.012 * 9.79\)   

    \(or, \rho = 902.2 Kg/m^3\)